# PainDiffusion: Can robot express pain?

### v1

@@ IV-C Qualitative results
Fig. 4 shows a comparison of a sample with the same stimuli and the corresponding expressions generated by the autoregressive baseline, ground truth, and PainDiffusion. It demonstrates that the expressions generated by PainDiffusion are more natural and closer to the ground truth, while the autoregressive baseline tends to maintain a nearly static facial behavior. In the comparison between diffusion forcing and full-sequence diffusion, qualitatively, diffusion forcing produces more stable predictions, whereas full-sequence diffusion introduces greater diversity and randomness, with less of a lead-follow relationship with the stimuli signal. We provide non-cherry-picked samples on the project website for better qualitative representation.
To evaluate the magnitude and area of facial movements in the generated sequences, we use the Farneback method [45] to detect motion. By varying the stimuli, emotion configurations, and expressiveness configurations, the mean movements are shown in Fig. 5. Stronger stimuli levels result in stronger facial motion. Neutral emotion status shows the least facial movement, while sadness and happiness lead to larger movements when responding to pain stimuli. Higher pain expressiveness tends to result in weaker movements.
Figure 5: Average movement observed under varying stimuli levels, emotion configurations, and expressiveness configurations, while keeping other metrics constant. Higher stimuli levels correspond to greater movement, though different emotions exhibit varying levels of movement. Additionally, increased pain expressiveness tends to result in weaker overall movement.
Figure 6: PSPI [38] sequence of the Autoregressive baseline and PainDiffusion across a sequence and their movement area.
@@ IV-D Quantitative results
TABLE I: Baselines. Comparison of our approach with the baseline.
PSPI [38] Expression Stimuli
PainSim ()($\downarrow$)( ↓ ) PainCorr ()($\uparrow$)( ↑ ) PainDist ()($\downarrow$)( ↓ ) PainDivrs ()($\uparrow$)( ↑ ) PainVar ()($\uparrow$)( ↑ ) PainAcc ()($\cdot$)( ⋅ )
Ground truth 0 0.9999 0.0000 0.0000 0.0552 17416
Random Training 342 0.5745 0.2273 0.1901 0.0163 17952
Nearest Neighbor 345 0.6773 0.2218 0.0000 0.0198 17864
FSQ-VAE Autoregressive [44] 297 0.3990 0.2338 0.0326 0.0159 17620
PainDiffusion w/ Full-seq Diffusion 216 0.4126 0.1578 0.0865 0.0454 18138
PainDiffusion w/ Diffusion Forcing 224 0.5839 0.1060 0.0638 0.0252 18110
()($\cdot$)( ⋅ ) means closer to ground truth is better.
Table I presents a comparison between our proposed method and other baseline approaches. Overall, our method outperforms the quantized-vector and autoregressive model across all metrics. Notably, our evaluation methods provide a more detailed perspective on the pain generation problem by using metrics based on the PSPI signal from both the generated output and the ground truth. The PSPI signal generated by PainDiffusion is more closely aligned with the ground truth compared to the heuristic and autoregressive baselines, as evidenced by lower PainSim and higher PainCorr.
In terms of expression diversity, PainDiffusion generates facial expressions that are not only closer to the ground truth but also exhibit greater variation across multiple predictions (lower PainDist, higher PainDivrs). Additionally, during a prediction sequence, our model achieves the highest expression variance, as indicated by higher PainVar. Despite this higher variance, analysis using the Farneback method [45] shows that PainDiffusion’s facial movements are more concentrated in key areas such as the mouth, eyebrows, chin, and nose (Fig. 6), which closely mirrors the ground truth (Fig. 4).
Regarding stimuli similarity, PainDiffusion achieves performance closer to the ground truth than the autoregressive baseline, with PainAcc being more aligned with the ground truth. As shown in Fig. 6, the stimuli signal tends to have a smoother shape, while the ground truth PSPI contains more fluctuations. It illustrates that PainDiffusion more closely follows the ground truth, particularly during periods of rapid stimuli change, indicating its sensitivity to dynamic pain variations. In contrast, autoregressive baseline shows less responsiveness to stimuli changes and deviates from the ground truth. Therefore, predictions that closely match the natural variability of the ground truth can be considered more accurate and realistic.
@@ IV-E Ablation
TABLE II: Ablation Study Results. Comparison of different configurations.
PainSim PainCorr PainDist PainDivrs PainVar PainAcc
Context Window 2 322 -0.0110 0.1513 0.0771 0.0284 18302
Context Window 4 323 -0.0137 0.1337 0.0736 0.0278 18282
Context Window 8 332 -0.0059 0.1075 0.0678 0.0304 18286
DF Uncertainty 0.5 304 0.0195 0.1900 0.0847 0.0440 18296
DF Uncertainty 1 314 0.0265 0.1788 0.0795 0.0394 18299
DF Uncertainty 2 319 0.0332 0.1383 0.0743 0.0304 18290
DF Uncertainty 4 368 0.0224 0.1269 0.0684 0.0283 18325
Guiding 1_1_1 * 333 0.0147 0.1172 0.0634 0.0228 18285
Guiding 1_2_4 * 316 0.0033 0.1144 0.0651 0.0220 18273
Guiding 0.5_1_2 * 316 -0.0005 0.1267 0.0713 0.0276 18272
Guiding 0.25_0.5_1 * 329 0.0169 0.1354 0.0740 0.0278 18291
* Guiding strengths respectively are emotion configuration, pain expressiveness configuration, and pain stimuli signal.
We analyze the effect of different context sizes, diffusion forcing uncertainty, and guiding strength on performance, with the results presented in Tab. II. A longer context window results in a smaller expression distance from the ground truth (PainDist) and a better correlation with the PSPI signal (PainCorr). It also leads to greater expression variation across predictions. However, this comes with the tradeoff of increased prediction time, as a longer context window shortens the prediction horizon.
Regarding diffusion forcing uncertainty, this parameter controls the level of uncertainty the model considers when generating new windows. Higher uncertainty requires more steps for generation. Our results show that an uncertainty level of 2 produces the best correlation between the generated PSPI signal and the ground truth (PainCorr). Lower uncertainty levels yield better diversity (PainDivrs, PainVar) but result in greater deviation from the ground truth (PainDist).
In terms of guiding strength, we find that prioritizing stimuli signals, followed by pain expressiveness and then emotion configuration, produces the best results. Weaker guiding strength increases diversity (PainDivrs and PainVar) but sacrifices expression similarity to the ground truth (PainDist).
The choice of configuration can vary based on whether the focus is on the accuracy of pain expression levels or the diversity of pain expressions. Therefore, selecting the best configuration depends on which aspect—correctness or diversity—takes precedence in a given application.
@@ V CONCLUSION
The current approach focuses solely on non-verbal expressions and limits controllability to emotional status and pain expressiveness. Future work could explore additional parameters, such as wear characteristics of materials or actual physical damage, to generate more accurate pain expressions. Incorporating auditory pain expressions, such as sounds, could also provide a more natural and multimodal way of expressing pain.

### v2

@@ IV-C Qualitative results
In the first phase of our qualitative evaluation, we analyzed the distribution of clinicians’ ratings from real-time rehabilitation exercise experiments, as shown in Figure 4. The results indicate that PainDiffusion generates relatively weak reactions, which aligns with the characteristics of the naturalistic dataset it was trained on. However, we acknowledge that clinicians expected reactions to align more closely with Japanese cultural norms, where expressions of pain tend to be more restrained. As anticipated, our model does not fully capture this cultural specificity, though it was judged to produce reactions that bear some resemblance to how Japanese patients typically express pain. Additionally, clinicians noted that the absence of a reference baseline for pain reactions made it difficult to assess whether the generated expressions were entirely appropriate. This highlights the need for calibration when deploying automated pain expression models in clinical applications to ensure cultural and contextual suitability.
In the second phase, the first video-based survey showed that PainDiffusion achieved a win rate of 31.2% ± 4.8% against the ground truth in terms of perceived realism, suggesting that PainDiffusion is capable of expressing pain in a way that is convincing to human observers. Since naturalistic pain expression datasets typically exhibit weaker reactions [6], and the BioVid HeatPain Database reflects this characteristic, PainDiffusion remains consistent with observations in the dataset. However, as illustrated in Figure 5, the model exhibits slightly weaker temporal consistency compared to the ground truth. We further evaluated diversity along three dimensions: movement amplitude, movement type (fast or slow), and overall variability. As shown in Figure 7, the model generates greater diversity in the eye region but lower diversity in the mouth region. Overall, it achieves moderate diversity across both movement size and type. These findings suggest that PainDiffusion is not only capable of generating realistic pain expressions but also has the potential to serve as a viable replacement for real patients in clinical training and simulation settings.
Figure 7: Diversity of generation of Paindiffusion assessed by clinicians.
We employed the Farneback method [39] to detect motion, evaluating the magnitude and area of facial movements in the generated sequences. To assess controllability, we varied the stimuli intensity, emotional configurations, and expressiveness settings. The mean movements are presented in Fig. 8. Our findings indicate that stronger stimuli levels elicit more pronounced facial motions. Neutral emotional states correspond to minimal facial movement, whereas sadness and happiness lead to more extensive movements in response to pain stimuli. Interestingly, higher expressiveness settings result in less intense movements.
Figure 8: Average facial movement generated by PainDiffusion under varying stimuli levels, emotion configurations, and expressiveness configurations, while keeping other configurations constant. Higher stimuli levels correspond to greater movement, though different emotions exhibit varying levels of movement. Additionally, increased pain expressiveness tends to result in weaker overall movement.
@@ IV-D Quantitative results
TABLE I: Baselines comparison.
Modalities PSPI [34] FLAME Params [23]
Metrics Sim $\downarrow$ Corr $\uparrow$ Dist $\downarrow$ Divrs $\uparrow$ Var $\uparrow$
($10^{-3}$)
Ground truth 0 999.9 0.00 0.00 0.06
Naive Methods
Random Training Sample $353^{\scriptscriptstyle\pm 139.4}$ $650^{\scriptscriptstyle\pm 5.0}$ 0.28 0.23 0.02
Nearest Neighbor $342^{\scriptscriptstyle\pm 0}$ $715^{\scriptscriptstyle\pm 0}$ $0.27$ $0.00$ $0.03$
Model-based Methods
FSQ-VAE Autoregressive [38] $299^{\scriptscriptstyle\pm 1.17}$ $396^{\scriptscriptstyle\pm 1.4}$ 0.23 0.03 0.02
PainDiffusion w/ Full-seq Diffusion $218^{\scriptscriptstyle\pm 1.85}$ $499^{\scriptscriptstyle\pm 2.6}$ $0.16$ $0.09$ $0.05$
PainDiffusion w/ Diffusion Forcing $226^{\scriptscriptstyle\pm 0.89}$ $597^{\scriptscriptstyle\pm 2.2}$ $0.10$ $0.06$ $0.02$
Table I presents a comparison between our proposed method and other baseline approaches. Overall, our method outperforms the autoregressive baseline across all metrics. Notably, our evaluation methods provide a more detailed perspective on the pain generation problem by using metrics based on the PSPI signal from both the generated output and the ground truth. As illustrated in Fig. 6, the PSPI signal generated by PainDiffusion is more closely aligned with the ground truth compared to the autoregressive baseline, and heuristic baselines as evidenced by lower PainSim and higher PainCorr in both of its variants.
Figure 9: Average of area of movement of a validation sample from PainDiffusion and Autoregressive baseline.
In terms of expression diversity, PainDiffusion generates facial expressions that are not only closer to the ground truth but also exhibit greater variation across multiple predictions (lower PainDist, higher PainDivrs). Additionally, during a prediction sequence, our model achieves the highest expression variance, as indicated by higher PainVar. Despite this higher variance, analysis using the Farneback method [39] shows that PainDiffusion’s facial movements are more concentrated in key areas such as the mouth, eyebrows, chin, and nose (Fig. 9), which closely mirrors the ground truth.
@@ IV-E Ablation
We conducted ablation studies to explore the hyperparameter space of PainDiffusion. To optimize computational resources, we evaluated a subset of the validation set to compute the relevant metrics. The results are summarized in Table II. Our findings suggest that the context window size significantly impacts both temporal coherence and diversity in the generated expressions. From experiments with a window size of 64 frames, we observed that a context window of 16 frames achieves the best balance between temporal consistency and diversity. Additionally, a diffusion forcing uncertainty value of 2 yielded the most stable results. For guidance strengths, we found that setting values of (0.5, 1.00, 2.00) for emotion configuration, pain expression configuration, and stimulus signal, respectively, provides an optimal trade-off between PSPI metrics and expression diversity metrics.
TABLE II: Ablation Study: Hyperparameters search.
Ablation Sim Corr Dist Divrs Var
$10^{-3}$ $10^{-1}$ $10^{-1}$ $10^{-1}$
Context Window Size
8 frames 305 639.1 1.16 0.68 0.26
16 frames 295 503.4 1.04 0.64 0.25
32 frames 301 529.9 0.89 0.59 0.25
Diffusion Forcing Uncertainty
0.5 314 422.0 1.57 0.84 0.44
1 307 505.5 1.40 0.74 0.35
2 279 624.8 0.99 0.63 0.25
4 306 496.4 0.95 0.62 0.25
Guiding Strength
Emo. Exp. Sti.
1.00 1.00 1.00 300 592.7 0.85 0.53 0.19
1.00 2.00 4.00 295 353.1 0.85 0.54 0.19
0.50 1.00 2.00 296 558.6 0.97 0.61 0.22
0.25 0.50 1.00 311 380.2 1.02 0.63 0.24
@@ V CONCLUSION
The current approach focuses exclusively on non-verbal expressions, with controllability limited to emotional state and pain expressiveness. Future work could extend this framework by incorporating additional factors, such as material wear characteristics or real-time physical damage, to enhance the accuracy of pain expression synthesis. Integrating auditory cues, such as vocal reactions and breathing patterns, could further improve realism by enabling a more natural and multimodal representation of pain. Additionally, the availability of a more diverse and naturalistic facial pain dataset would be beneficial in improving the generalization of automatic synthesis methods across varied demographics and clinical contexts.